OpenStudy (anonymous):

Find the distance between the given points R(0, 0) and S(6, 8) Distance= √(22) 10 2√7

OpenStudy (anonymous):

10

OpenStudy (anonymous):

a^2+b^2=c^2 (6^2)+(8^2)=100 =sqrt{100} =10